OpenStudy (photon336):

If the boiling point constant of water is 0.5 C/m what is the approximate boiling point of 2m solution of magnesium chloride? Answers in Celsius 101.5 98.5 97.0 103.0

OpenStudy (photon336):

\[MgCl _{2} --> Mg ^{+} + 2Cl ^{-}\] one mole of Magnesium ions 2 moles of chlorine ions. 1+2 = 3 moles of ions i = 3. m = molality moles/kg solvent. 2 \[iK _{b} \Delta T \] \[3*0.5\frac{ C}{ m }*2m = 3C = \Delta T \]

OpenStudy (photon336):

\[\Delta T = t _{f}-t _{i} \] boiling point is 100C so we know that the boiling point had to have been raised over that by some amount. \[\Delta T + Ti = Tf \] \[3C + 100 = 103 C\]